OpenStudy (anonymous):

calculate the number of cations in 7.45g of lithium sulfate

OpenStudy (anonymous):

n= m/M n=N/L -> N = n*L Li2SO4 -> 2(Li^+) + (SO4^2-) N(Li+) = 2* N(Li2SO4)